The wind pushes a paper cup along the sand at a beach. The cup has a mass of 25kg and accelerates at a rate of 5m/s2. How much force is the wind exerting on the cup?
1
Expert's answer
2020-06-26T14:37:35-0400

By the Newton's Second Law of Motion, we get:


"F = ma = 25 \ kg \cdot 5 \ \dfrac{m}{s^2} = 125 \ N."

Answer:

"F = 125 \ N."